Ramirez Corporation sells two types of computer hard drives. The sales mix is 30% (Q-Drive) and 70% (Q-Drive Plus). Q-Drive has variable costs per unit of $90 and a selling price of $150. Q-Drive Plus has variable costs per unit of $105 and a selling price of $195. Ramirez’s fixed costs are $891,000. What is the total break-even point?

  

3,300

   

4,455

   

11,000

   

7,700

NOTE: IT IS NOT 3300 APPARENTLY ACCORDING TO THE PROGRAM

(If it is 3300, let me know so that i can speak to an administrator to mark my answer as correct)
Answer #1

Answer:

Option 3rd is correct. i.e 11,000 units

Contribution margin=Selling Price per unit−Variable Cost per unit
(Q-Drive)              = $150-90 = $60
(Q-Drive Plus)    = $195-105 = $90
Combined contribution margin= ($60* 30%)+($90*70%) = $81
Total Breakeven sales = Fixed cost/Contribution margin per unit(total)
                                = $891,000/$81 per unit
                               = 11,000 units
Answer #2

3300 is the answer